OpenStudy (anonymous):

Medal and Fan How many solutions does the equation have? 3(4x -2) = 2(6x - 3) A. One Solution B. Infinite Solutions C. No Solution

OpenStudy (anonymous):

3(4x-2)=2(6x-3) 12x-6=12x-6

OpenStudy (anonymous):

So one solution

OpenStudy (anonymous):

since both side are the same you can get infinity solutions

OpenStudy (anonymous):

put in a random number and see

OpenStudy (anonymous):

oh ok thx

OpenStudy (anonymous):

anytime